### v1

# I Wish to Have an Argument: Argumentative Reasoning in Large Language Models

Abstract
We evaluate the ability of contemporary large language models (LLMs) to perform argumentative reasoning. We frame our experiments in terms of the argument mining (AM) and argument pair extraction (APE) tasks, and evaluate their ability to perform reasoning at increasing levels of abstraction in the input and output representations (e.g., arbitrary label sets, semantic graphs). We find that, although LLMs are able to match or surpass the state-of-the-art in AM and APE, their argumentative reasoning performance is very dependent on the input and output representation. We also find an “exemplar effect”, where too many exemplars increasingly become detrimental for task performance, and about $4-5$ being the optimal amount. Neither result extends to chain-of-thought (CoT) prompting: we find the exemplar effect to be nullified, and our results suggest that CoT allows for better performance under ill-conditioned problems. We hope that the work reported contributes to the improvement of argumentative reasoning in LLMs.
1 Introduction
Large language models (LLMs) such as GPT-4 (Open AI, 2023) are able to solve a significant amount of natural language processing tasks to a spectacular degree of accuracy. Lately, there has been a push to use these models to automate more complex tasks reliant on deeper comprehension of language, such as data labelling (Cheng, Li, and Bing, 2023; Liu et al., 2023), model evaluation (Zheng et al., 2023), and even paper reviews (Liu and Shah, 2023).
Argumentative reasoning is central to all of these efforts. This type of reasoning encompasses formal and informal logic, and has as a prerequisite a deep understanding of language and the context. Plenty of work has been carried out to evaluate LLM performance in formal logic scenarios; but, in human-centered applications, such as chatbots, formal logic issues are overlooked in favor of defeasible reasoning (Walton, 2008). Hence, to accurately decide whether LLMs are reliable on more complex tasks, such as data annotation and human interaction, we must also evaluate their capabilities in the context of argumentative reasoning.
In this paper we perform one such evaluation for GPT-3 (Brown et al., 2020) and GPT-4. We explore these model’s argumentative reasoning capabilities by altering the input and output representations and subsequently observing their performance in two argumentation-related tasks. We frame our experiments from an argument mining (AM) and argument pair extraction (APE; Cheng et al., 2020) perspective. AM requires the model to be able to recognize spans of a text as belonging to an argument; and APE tasks the model to match arguments from two texts as argument-rebuttal pairs.
For the input and output representation evaluation, we test progressively more abstract reasoning settings: from the concrete setting, where the model is tasked to return matches from a passage (e.g., in AM it would return the spans that correspond to an argument); to the purely symbolic setting, where we require from the model to return the answer with arbitrary symbols (e.g., extract the spans, label them based on predefined rules, and only return the labels), or the input is in an abstract format such as abstract meaning representation (AMR; (Banarescu et al., 2013)) graphs. AMR graphs are known to aid in argument comparison and interpretation (Opitz et al., 2021).
We also factor in the number of exemplars to test if the model’s reasoning capabilities depend on their context length, and chain-of-thought (CoT; (Wei et al., 2022b)) prompting, a prompting paradigm that conditions the model to work step-by-step by following a templatized process (“Let’s think about this step-by-step…”). It is known to provide good results in multiple reasoning tasks (Wei et al., 2022b; Kojima et al., 2022; Suzgun et al., 2023). The diagram below displays our evaluation.
[…]
1.1 Findings
In terms of raw scoring, we find that GPT-4 is able to reach SOTA performance in APE, and near-SOTA in AM. However, our analysis also shows the following:
  1. 1.
    Both LLMs fall short when the input and output representations are not carefully conditioned. Concrete reasoning does well, but low-abstraction (i.e., with “hints” as the input representation) symbolic approaches outperform it.
  2. 2.
    The performance with respect to the number of exemplars behaves consistently regardless of setting, representation, or task. The LLMs have best performance when observing between 4 and 5 exemplars. More exemplars (up to 40 in some cases) cause a linear decrease in performance.
  3. 3.
    Performance drops noticeably, but remains above random, when the input representations are too abstract (e.g., AMR graphs). This does not hold for purely symbolic output representations (e.g., matrices), where the performance drops to near-zero levels.
  4. 4.
    The results above do not extend to CoT. We find that, in line with the literature, CoT outperforms analogous non-CoT approaches. But we find that the distributions output by CoT are similar, which suggests that the effect of the input and output representations changes is mitigated by this technique. The “exemplar effect” mentioned above does not extend to CoT either.
2 Methodology
2.1 Data
Throughout this paper, we utilize the Review-Rebuttal Submission-v2 (RRv2) dataset from Cheng et al. (2020). It is comprised of $4,764$ pairs of review and rebuttal passages related to scientific article submissions. The test set size is $474$ pairs, and the data was human annotated by $5$ annotators following a specific guideline. Each passage is sentence-separated, and normally includes multiple arguments. RRv2 is a comprehensive corpus and focuses on long-distance relationships between statements; in addition to having two tasks using the same data (AM and APE), which are both actively researched in argumentative reasoning.
For AM, each sentence is labelled with a BIO scheme, and the model must retrieve (label) each sentence from the review and rebuttal entries. [^1] In the RRv2 corpus, the BIO tags correspond to the Beginning, Inner, and Outer parts of an argument. In this task, the distinction between review and rebuttal is not relevant and each entry is treated as a separate point in the corpus.
For APE, the arguments within each review-rebuttal pair must now be aligned: every argument made by a reviewer must be mapped, when applicable, to a refutation from the rebuttal. This is normally represented as a binary matrix with overlaps (Bao et al., 2022; Cheng et al., 2021). See Figure 1 for examples of the tasks and its labelling scheme. Prior to use we clean the text from tag and sentence delimiters.
2.2 Models Evaluated
We evaluate GPT-4 and the text-davinci-003 variant of GPT-3. The latter is similar to the original GPT-3 model, but further pretrained with reinforcement learning with human feedback (Ouyang et al., 2022; Christiano et al., 2017) and instruction pretraining (Wei et al., 2022a; Sanh et al., 2022; Ouyang et al., 2022). It is sometimes referred to as GPT-3.51, and has a context length of $4,097$ tokens. [^2] https://platform.openai.com/docs/models
Figure 1: Examples of the (raw) data entries in RRv2. The BIO tag is in parentheses, and highlighted in orange and blue the “B” and “I” labels corresponding to an argument. In AM, review and rebuttal (response) passages are considered independent, and the task is to label all the arguments from each passage. In APE, the task is to match arguments from the review to their corresponding refutations. In the above the first argument is unmatched, and the second (“BERT produces…”) pairs to the first argument from the response.
Few details have been released around the architecture, model size, or pretraining procedure for GPT-4. Its authors consider it marginally better than GPT-3 at some tasks, and better at more complex tasks. They evaluate it in benchmarks specific to GPT-4 that do not explicitly include AM or APE. However, some of these benchmarks have an argumentation component, such as the essay portion of the SAT. [^3] A test for entrance to universities in the United States. The variant of GPT-4 we utilized has a context length of $32,768$ tokens. Throughout our experiments, for both models we set the temperature to $0.8$, set the maximum return tokens based on the task, and left all other parameters as their default. To account for randomness, we report the average of five calls.
2.3 Prompting
The models are well-known to be sensitive to the prompt’s wording (Lu et al., 2022), so prior to starting the work carried out some small experimentation to find out which phrasing would be best. We always include exemplars of the desired output format, and, when the prompts used $n$ exemplars, we used the first $n$ points from the development set. In the case of AM, we also prompted GPT-4 with and without CoT, as it has been shown to improve model’s reasoning capabilities in various tasks. We wrote all our prompts following the structure from Open AI (2023), as we observed it produced more reliable outputs. For CoT we followed the standard phrasing (“Let’s read line-by-line and solve this step by step”) and indicated which sentence was being read, as well as the rationale. For example, “We now read {SENTENCE}. It follows the previous argument, and hence it is labelled with an ‘I’.” Sample prompts can be found in Figures LABEL:fig:sampleamconcrete and 2.444The full set of the prompts, scoring code, and model responses will be available at https://github.com/adewynter/argumentation-llms We used the Open AI API for all our experiments, and a consumer-grade laptop for the data processing.
2.4 Scoring
[…]
2.5 Baselines
[…]
Extract from the passage all the arguments. Give
every argument in its own paragraph.
The output should be in the form:
|begin response|
|START| line from argument 1
line from argument 1
|START| line from argument 2
line from argument 2
etc
|end response|
For example,
{EXEMPLARS}
Passage:
{PASSAGE}
Response:
[…]
    Exemplars, if any, are included in the prompt and follow the output representation.
    Otherwise (in the case of zero-shot), we only leave the specified output representation. More complex prompts, such as the one for APE, simply ask for the pairing (e.g., return all arguments from the response that match those of the review”).
[…]
Extract from the passage all the arguments.
Label the beginning of every argument with a B”.
Label the rest of the argument with an I”.
Label every line that is not part of an argument
with an O”.
The output should be in the form:
|begin response|
B
I
O
etc
|end response|
For example,
{EXEMPLARS}
Passage:
{PASSAGE}
Response:
|begin response|
Figure 2: Sample symbolic AM prompt. The model is requested to parse each line and return BIO tags. Exemplars, if any, are included in the prompt, and follow the specified output representation. For zero-shot, we only specify the output representation. In other output representations the model is requested to mark the B-label (e.g, “(15) 16 17”). In APE we request the index-based response with matching arguments (e.g., “15, 16, 17: 9, 10, 18, 19”) and convert it into the binary matrix representation for scoring. We also experimented with requesting the binary matrix outright.
2.6 Settings
We have named our settings (i.e., representations) as concrete and symbolic, to broadly distinguish the approach taken towards representing the task. The former returns full sentences, while the later encompasses a variety of input and output symbols. This is solely for practical purposes: in reality, symbolic approaches cover a variety of abstractions which may be considered “easier” than the concrete approach, and, strictly speaking, the concrete approach is a symbolic approach (Brachman and Levesque, 2004). Below we describe these settings, and we discuss our distinctions further in 3.3.
2.6.1 Concrete
We instruct the model to return a portion of the text based on the instructions given: in AM, the lines corresponding to an argument, and in APE, the argument pairs themselves. To distinguish the “B” and “I” labels, we enforce a specific return format to work with our parsing code via a special token (|START|). For scoring concrete answers, we always expect an exact match. See Figure LABEL:fig:sampleamconcrete for a basic AM concrete prompt.
2.6.2 Symbolic
The model is instructed to return symbols (labels) based on its judgment of the text. This requires more reasoning steps than the concrete setting, given that the model is solving AM and labelling the span with a pre-specified, arbitrary label set from the instructions. See Figure 2 for an example. In the AM symbolic setting we evaluated two types of labels: the BIO tags and the line indices. For APE we evaluated the line indices and the full matrix representation. For AMR, we used amrlib [^5] https://github.com/bjascob/amrlib and a pre-trained BART-based model (Lewis et al., 2020) included in the library. The AMR graphs were created at the sentence-level.
3 Experiments and Results
3.1 AM/APE: Symbolic and Concrete Reasoning
In our first experiment we evaluated the performance of the models in AM and APE in both the concrete setting and various iterations of symbolic settings. The results for the best-performing prompts and settings are in Table 1, and a description of every symbolic setting in Table 2.
[…]
In AM, the best-performing symbolic setting had line indices included in the input representation, and requested the indices of each argument as the output representation. To handle the conversion to a BIO label set, we instructed the model to return the “B” labels as indices enclosed in parentheses. Not including the indices in the input did not lead to an equivalently good performance.
In APE, requesting a full binary matrix output led to extremely poor performance ($9.88\%F_{01}$, well below the random baseline, and closer to guessing all zeros every time). However, both symbolic and concrete approaches performed well with respect to the existing non-LLM-based baselines. We tested other approaches, such as first extracting the arguments and then matching them, but it did not yield sufficiently good results. Due to token-length and budget limitations, we were unable to test CoT and AMR in APE.
CoT approaches had generally better performance than their non-CoT counterparts, even though they use fewer exemplars due to their higher token usage. The only exception to this observation was the best-performing approach (AM: indices inline and indices as output) where the difference was of $3\%$ points–still above the concrete setting.
[…]
GPT-3 (c) $39.86\pm 0.51$ $18.58\pm 0.70$
GPT-4 (c) $64.51\pm 0.53$ $53.84\pm 0.73$
GPT-3 (s) $62.00\pm 0.32$ $20.15\pm 0.91$
GPT-4 (s) $70.63\pm 0.21$ $49.85\pm 0.96$
MRC-APE $72.43$ $39.92$
MLMC $71.35$ $32.81$
Table 1: Results for the AM and APE tasks in the symbolic (s) and concrete (c) settings. We report the best-performing prompt for an average of five calls: in the case of symbolic reasoning, the best performing approach in both tasks had indices inline and indices as the output. We also report MLMC and MRC-APE, the two best-performing, non-LLM-based approaches for RRv2. Overall, GPT-4 had good performance, almost matching the existing baselines in AM and outperforming them in APE. GPT-3 had much lower performance.
3.2 Performance and Number of Exemplars
On our second experiment we evaluated the performance of the LLMs in AM and APE, but compared with the number of exemplars seen by the LLMs. The number of exemplars was $\{0,4,8,16,\tau\}$, where $\tau$ is the maximum number allowed by the model’s token limit. In the case of GPT-3, $\tau$ tended to be around $4$; for GPT-4 it was a variable number depending on the task and setting, with an average of $44$ for symbolic, non-CoT approaches in AM ($15$ CoT, $9$ AMR) and $29$ concrete; and $22$ for both symbolic and concrete in APE. The total size of our collected data, including all exemplars, tasks, settings and five calls, is about 3.4GB. The results are in Figures 3 and 4. The models showed on average a peak in performance when the number of exemplars was $4$, and their performance steadily decreased as this number increased. This peak was independent of the task and the setting, since all curves were similar to one another. In the case of CoT, we did not observe such a trend.
Figure 3: Effect of the number of exemplars in AM. Thick blue line is the concrete approach, thickest red line is the average symbolic performance. Last entry in all plots is the average maximum number of exemplars supported. On average, the LLMs performance peaked at $4$ exemplars, and steadily decreased as the exemplars became more numerous. The exception to this observation was CoT prompting (solid red lines), where they consistently outperformed concrete prompting approaches and all but one of the symbolic approaches (dashed red lines), and did not show this decreasing trend. GPT-3’s performance was quickly outclassed by GPT-4, even under the same number of exemplars.
Figure 4: Effect of the number of exemplars in APE. Thick blue line is the concrete approach, thickest red line is the average symbolic performance. Last entry on each plot is the average maximum number of exemplars supported. Similar to AM, on average model performance peaked at $4$ exemplars, and steadily decreased as they became more numerous. We were unable to test CoT and AMR under this approach due to limitations on token length and budget. Dotted lines for symbolic are the various symbolic representations.
3.3 Performance and Input Representation
On our last experiment we explored whether the representation of the input text into progressively more abstract ways led to a change of performance in the symbolic performance of the models. In this section we focused solely on GPT-4 and AM, since the AMR graphs tended to be too big for the models to handle (for GPT-3 in both AM and APE, for GPT-4 only APE).
We tested three representations (text with indices inline, plain text, and with and without an AMR graph). There is no rigorous way to quantify the level of abstraction for these approaches. However, we consider the concrete approach to be least abstract, “hints” (indices inline, indices in output) to be slightly more abstract, and purely symbolic input representations (AMR graphs) as the most abstract. In between these categories, we ranked them based on the complexity of the output representation: BIO tags are more abstract than indices, since they have rules for matching, and both are more complex than concrete approaches since outputting the matching sentence is simpler than mapping to an arbitrary symbol and returning it. The full list of experiments is in Table 2, and our results in Figure 5.
In non-CoT approaches, we observed noticeable improvements when the abstraction level was sufficiently low (indices inline and indices in output; concrete), and as the abstraction increased, the model started to perform poorly, but retained above-random performance. These results are significant under a Welch’s $t$-test carried out on the predictions arrays: when the input text is unaltered, requesting line numbers or BIO tags as the output provides no difference in predictions with CoT ($p\approx 0.77$; large $p$-values imply the distributions have identical expected values), but does provide a noticeable difference without ($p<0.05$).
[…]
Task Input Output (label set)
AM* Text Text and |START|
AM* Text with indices Indices
AM* Text Indices
AM* Text BIO tags
AM Text with AMR graph BIO tags
AM AMR graph BIO tags
Table 2: Input representations tested, in roughly increasing order of abstraction. Tasks marked with an asterisk (*) were tested with and without CoT. The first row (input text, output text requested) corresponds to our concrete reasoning setting. Our ranking of abstraction is arbitrary: we consider the text with indices marked inline less abstract than a text without them, since the former provides a “hint” of what the label set is supposed to be like. Output representations with BIO tags as the output are more abstract, since it requires some rule matching in order to determine which symbol corresponds to which part of the problem. AMR graphs are considered the most abstract.
Figure 5: Effect of abstraction level in the input/output representation with respect to output performance in AM. For non-CoT approaches, the more abstract (under our criterion) the input representation, the more difficult it is for the model to solve the task. For example, if the input is conditioned to have “hints” about the desired output (for example, requesting indices and adding line numbers to the input text) the performance increases significantly. That said, even the most abstract representations (AMR graphs) outperform noticeably the random baseline. CoT approaches (at the top, marked with an asterisk) consistently outperformed their non-CoT counterparts, even when the maximum number of exemplars supported is much lower, and regardless of abstraction level. We did not test CoT with AMR in AM due to token limitations.
4 Discussion
4.1 Relationship of F-scores and Reasoning
The models, and GPT-4 in particular, performed well in both AM and APE, to the point of beating, or almost matching, the existing SOTA models. This alone is not sufficient to claim that the models are able to perform argumentative reasoning, let alone understand it.
When altering the input and output representation with conceptually minor changes–such as adding line numbers or requesting a BIO set as opposed to integers–the models had noticeably different performances. We argue that this is indicative of a potential false positive: the models are unable to reason in an argumentative setting, but give excellent appearance of being able to do so.
It is worth noting that, although GPT-4’s performance in APE is SOTA, it is still well-below the equivalent performance in AM. Although the tasks are slightly different, another plausible explanation is that the models failed to generalize due to the length of the task, which it is known to cause problems in transformer-based models (Anil et al., 2022).
4.2 Impact of CoT in Argumentative Reasoning
That said, we noticed that this performance difference did not extend to the inclusion of CoT. These prompts had on average better performance than their non-CoT counterparts, and we observed that it yielded better results in ill-posted (here, overly abstract) problems. However, our analysis showed that the output distributions for all CoT approaches were rather similar. This is perhaps indicative that CoT allows the models to return the desired output regardless of representation.
One possible explanation for the similar output distributions is due to the templatized nature of CoT: it usually starts with a common phrase (in our case, “Let’s read line-by-line and solve this step by step”) and follows very specific steps that represent (in our prompt) the reasoning required. While we don’t discard this possible explanation, it is worth noting that the input and output representations are completely different in all approaches. That said, it is clear that CoT mitigates the effect of input and output representations on the execution of the task.
4.3 The Exemplar Effect
Perhaps the most surprising aspect of our evaluation was the effect of the number of exemplars with respect to the model’s downstream performance. It was not a linear relationship, and instead there was a clear “sweet spot” at around $4$ samples. Originally, Brown et al. (2020) had indicated that more exemplars in GPT-3 would yield better performance, but that this number must be finetuned ahead of time. We observed the same thing, but these results did not hold for GPT-4. What this suggests, from our experiment results, is that (assuming that more exemplars imply better understanding of the task) the models are not accurately performing inference from the data provided.
That said, this “exemplar effect” did not extend to CoT settings, since the maximum number of tokens used had slightly higher performance than the other approaches. However, we do not discard the possibility that models with longer token limitations could also show a similar exemplar effect in CoT.
5 Limitations
5.1 Reasoning Evaluation
It could be argued that our results are not complete in terms of evaluating argumentative reasoning capabilities. We agree: recognizing an argument is not the same as deciding whether it is sound, valid, or neither. However, without the ability of the model to show that it is able to recognize arguments and identify relations between them, any potentially generated argument or result evaluating the model’s performance in these tasks could be deemed as untrustworthy. Note that issues arising from the model not understanding the prompt (Webson and Pavlick, 2022) are not present in our evaluation, since the results reported here all perform better than the random baselines, and our analysis was comparative in nature.
5.2 Data contamination
We factored out, to an extent, potential data contamination, which is known to impact downstream model performance (Carlini et al., 2023; Liang et al., 2022; Lee et al., 2023; De Wynter et al., 2023), by tasking the model to recognize arguments from the passage. However, we are unable to guarantee that the models have not been trained with this data, and therefore have at least some bias towards these results.
5.3 Model Setup
We only evaluated two models, so our results may not extend to other LLMs. We did not fine-tune the models, and opted instead to treating them as generalists performing in-context learning, in line with their standard, contemporary usage. However, it is likely that fine-tuning or specialized pre-training could show more insights into our results, and probably a more uniform performance distribution across representations: after all, MRC-APE and MLMC are specialist models and remain the SOTA in AM for RRv2.
6 Related Work
6.1 Informal Reasoning and Argumentation
In this section we limit ourselves to discuss informal reasoning and argumentation evaluation of LLMs. For an introduction to AM covering other modelling approaches see the survey by Lawrence and Reed (2020) and for a broader introduction to the field and open problems see Lauscher et al. (2022).
There has been work on using LLMs for AM (Zhao et al., 2023), as well as determining argument quality (van der Meer et al., 2022; Holtermann, Lauscher, and Ponzetto, 2022). More recently, Han et al. (2022) showed that GPT-3 is unable to mimic human-like inductive reasoning. While non-LLM-based approaches to APE do exist (Cheng et al., 2020; Bao et al., 2022), our work is the first to evaluate LLMs for this particular task.
A subset of informal reasoning, commonsense reasoning, is a very active of research. It is not strictly the same as argumentative reasoning, but it allows for the evaluation of some types of defeasible arguments. While not the direct subject of this work, there are reports that GPT-3’s performance is rather low in commonsense reasoning (Cong, 2022) and figurative language (Chakrabarty, Choi, and Shwartz, 2022). In addition, there are techniques to improve the model’s performance in this area, such as generating explanations for out-of-distribution observations (Brahman et al., 2021), corpora and benchmarks (Forbes et al., 2020; Rudinger et al., 2020; Pyatkin et al., 2023), and decoding (Cohen and Mooney, 2023) and prompting (Liu et al., 2022; Jin et al., 2022; Štefánik and Kadlčík, 2023; Lu et al., 2022) strategies.
Of note, Webson and Pavlick (2022) explored whether GPT-3 and other LLMs were able to understand the prompts. They found evidence to the contrary, and although they provided two alternate interpretations for their findings (the task was too hard or the model refused to answer), their main conclusion is that the models do not always follow, or understand, the instructions the way a human being would.
6.2 Formal Reasoning
The scientific community has long been interested in the evaluation of reasoning capabilities for statistical models, and LLMs are no exception ((Zhang et al., 2023); (Dziri et al., 2023); (Dan, Bastani, and Roth, 2022); (Liu et al., 2022); to name a few). Benchmarks like BIG-Bench Hard (Suzgun et al., 2023) have subsets of data that are meant to evaluate formal reasoning. Saparov and He (2023) performed a formal analysis of the reasoning capabilities of LLMs. They built an ontology and tested whether GPT-3 was able to perform modus ponens given a specific premise. They found that the models are unable to do so consistently, with a much lower performance under fictional worlds. This has been taken beyond formal logic and into real-world reasoning problems (e.g., maths problems), applying GPT-3, and other LLMs with much success (Open AI, 2023; Lewkowycz et al., 2022; Zong and Krishnamachari, 2023). That said, there exists evidence to the contrary on some tasks, such as arithmetic (Dziri et al., 2023). The latter reference also indicates that there is strong evidence that tasks with significant reasoning depths will cause these models to fail, a result also shared by Anil et al. (2022).
7 Conclusion
In this paper we evaluated the argumentative reasoning capabilities of GPT-3 and GPT-4. We altered the input and output representations of the data to try to understand their behavior via their performances in AM and APE. We found that the models present good performance in these tasks, but this was not sufficient to conclude that they were indeed performing any type of reasoning. On the contrary, our experiments showed that the input and output representations had significant impact on their downstream performance. This sensitivity suggested that the model’s applicability to critical uses such as data labelling and paper reviewing needs to be exercised with care.
However, when the input and output representations included CoT, our results became more surprising: the “exemplar effect” (too many exemplars being detrimental for performance) was no longer present, and the output distributions were very similar regardless of input and output representations. We were unable to conclude that CoT helped argumentative reasoning in LLMs, but we did note that it helped mitigate potential issues associated with ill-conditioned problems.
As mentioned in the Limitations section, our work may be expanded by exploring the ability of the models to working in-depth with argumentation. Our work did not evaluate their ability to judge an argument’s strength, or for the model to provide reasonable rebuttals. Moreover, due to token length limitations, we were unable to evaluate AMR, the most abstract setting we tested, with CoT. We believe that the performance difference between non-CoT AMR and CoT AMR would provide valuable insights on to what extent these models are able to discern abstract input representations successfully. All of these are well-known problems to the computational argumentation field and the scientific community at large. We plan to explore in further work these as they relate to LLMs (GPT-4 specifically, given its performance compared to GPT-3) as in-depth evaluation of these is important for downstream applications.
Finally, although the goal of our work is to study LLM argumentative reasoning capabilities, we believe that this type of representation-based analysis could be extended to other models and non-argument-related tasks, such as natural language inference.
Acknowledgements
The authors wish to thank Liying Cheng for answering questions about the RRv2 dataset and its methodologies.
Appendix A Full Prompts
In this section we show samples of our AM and APE prompts. [^6] Full prompts in https://github.com/adewynter/argumentation. For AM we include our concrete (Figure LABEL:fig:amconcrete), symbolic (BIO and indices inline; Figures LABEL:fig:amsymbolicbio and LABEL:fig:amsymbolicindices), and one CoT (BIO; Figure LABEL:fig:amsymbolicbiocot) prompts. For APE we include the concrete (Figure 7) and one symbolic (indices as output; Figure 6) prompts. Due to space, we have omitted sample inputs and part of the exemplars.
[…]
Here is a |review| and a |response|. Return all pairs
of argument and rebuttal from the text.
Arguments must come from the |review|.
Rebuttals must come from the |response|.
Example:
|start of review|
{EXEMPLAR REVIEW 1 GOES HERE}
 |end of review|
|start of response|
{EXEMPLAR RESPONSE 1 GOES HERE}
|end of response|
Answer:
|start of answer|
- argument 0 (5 6) : 1 2 3 4 5 6 7 8 9 10 11 12 13
|end of answer|
Passages begin:
|start of review|
{REVIEW}
 |end of review|
|start of response|
{RESPONSE}
|end of response|
Answer:
|start of answer|
Figure 6: APE symbolic prompt with one exemplar (omitted). There is only one match as a possible solution (lines $5,6$ in the review; lines $1-13$ in the response). In zero-shot scenarios we specify the desired output representation.
[…]
Here is a passage.
Return every line from the passage that is part of an
argument, and ONLY these lines. Mark the beginning of
every argument with |START|.
For example:
|passage start|
- The author studies the quantization strategy of
CNNs in terms of Pareto Efficiency.
- Through a series of experiments with three
standard CNN models (ResNet, VGG11, MobileNetV2),
the authors demonstrated that lower precision value
can be better than high precision values in term of
Pareto efficiency under the iso-model size scenario.
- They also study cases with and without depth-wise
convolution, and propose a new quantization method,
DualPrecision.
- Comment:
- I am not at all familiar with quantization methods,
therefore no knowledge of relevant related works.
- If, however, the authors did a thorough job of
surveying related works and chose sensible baselines,
I think the experiments demonstrate the usefulness
of the new DualPrecision technique.
|passage end|
Response:
- |START| I am not at all familiar with quantization
methods, therefore no knowledge of relevant related
works.
- If, however, the authors did a thorough job of
surveying related works and chose sensible baselines,
I think the experiments demonstrate the usefulness of
the new DualPrecision technique.
Passage:
|passage start|
{PASSAGE}
|passage end|
Response:
[…]
Here is a passage.
For every line in the passage:
- Mark the beginning of every argument with B.
- Mark the inside of every argument with I.
- Mark everything else with O.
For example:
|passage start|
{EXEMPLAR GOES HERE}
|passage end|
Response:
Lets think step-by-step and read it line-by-line.
We read: The author studies the quantization
strategy of CNNs in terms of Pareto Efficiency.”:
This sentence is not related to an argument, so we
mark it as O”.
Next we read: If, however, the authors did a
thorough job of surveying related works and chose
sensible baselines, I think the experiments
demonstrate the usefulness of the new DualPrecision
technique.”: This follows the previous sentence. So
we are still reading an argument. We mark it as
I”.
So the answer should be:
 - O
 - O
 - O
 - O
 - O
 - B
 - I
Passage:
|passage start|
{PASSAGE}
|passage end|
Response:
Lets think step-by-step and read it line-by-line.
We read:
[…]
        This prompt performs step-by-step reasoning on AM by following a templatized generation and specific labelling logic.
[…]
Here is a passage.
For every line in the passage:
- Mark the beginning of every argument with B.
- Mark the inside of every argument with I.
- Mark everything else with O.
For example:
|passage start|
|passage start|
- The author studies the quantization strategy of
CNNs in terms of Pareto Efficiency.
- Through a series of experiments with three
standard CNN models (ResNet, VGG11, MobileNetV2),
the authors demonstrated that lower precision value
can be better than high precision values in term of
Pareto efficiency under the iso-model size scenario.
- They also study cases with and without depth-wise
convolution, and propose a new quantization method,
DualPrecision.
- DualPrecision empirically outperformed 8-bit
quantization and flexible quantization methods on
ImageNet.
- Comment:
- I am not at all familiar with quantization methods,
therefore no knowledge of relevant related works.
- If, however, the authors did a thorough job of
surveying related works and chose sensible baselines,
I think the experiments demonstrate the usefulness
of the new DualPrecision technique.
|passage end|
Response:
 - O
 - O
 - O
 - O
 - O
 - B
 - I
Passage:
|passage start|
{PASSAGE}
|passage end|
Response:
[…]
Here is a |review| and a |response|.
Return all pairs of argument and rebuttal from the
text.
Arguments must come from the |review|.
Rebuttals must come from the |response|.
Example:
|start of review|
{EXEMPLAR REVIEW 1 GOES HERE}
 |end of review|
|start of response|
{EXEMPLAR RESPONSE 1 GOES HERE}
|end of response|
Answer:
|start of answer|
- argument 0
 - I am not at all familiar with quantization methods
therefore no knowledge of relevant related works.
  - If, however, the authors did a thorough job of
 surveying related works and chose sensible baselines
 I think the experiments demonstrate the usefulness
 of the new DualPrecision technique.
- rebuttal 0
 - We have updated our manuscript so that it states
 our contributions explicitly.
  - While our proposed DualPrecision is simple yet
  effective, we would like to emphasize that it is
  only one of our contributions.
  - For the reviewers convenience, we re-iterate our
  main contributions below:
  - Overall, we systematically analyze the model size
  and accuracy trade-offs considering both weight
  precision values and the number of channels for
  various modern networks architectures (variants of
  ResNet, VGG, and MobileNet) and datasets (CIFAR
  and ImageNet) and have the following non-trivial
  and novel contributions:
|end of answer|
Passages begin:
|start of review|
{REVIEW}
|end of review|
|start of response|
{RESPONSE}
|end of response|
Answer:
|start of answer|
Figure 7: APE concrete prompt (non-zero shot, but the exemplars have been omitted for brevity). The response is given in Markdown format, with the indentation denoting the interior parts of an argument. We also use an indexing system (“argument 0”, “rebuttal 0”) for matching.
[…]
Here is a passage.
For every line in the passage:
- Group and return the index of lines corresponding
to an argument.
- Put the index of the beginning of the argument in
parenthesis, and then list below the remaining
indices, if any.
- Do not return anything else.
For example:
|passage start|
- 0) The author studies the quantization strategy of
CNNs in terms of Pareto Efficiency.
- 1) Through a series of experiments with three
standard CNN models (ResNet, VGG11, MobileNetV2),
the authors demonstrated that lower precision value
can be better than high precision values in term of
Pareto efficiency under the iso-model size scenario.
- 2) They also study cases with and without depth-
wise convolution, and propose a new quantization
method, DualPrecision.
- 3) DualPrecision empirically outperformed 8-bit
quantization and flexible quantization methods on
ImageNet.
- 4) Comment:
- 5) I am not at all familiar with quantization methods,
therefore no knowledge of relevant related works.
- 6) If, however, the authors did a thorough job of
surveying related works and chose sensible baselines,
I think the experiments demonstrate the usefulness of
the new DualPrecision technique.
|passage end|
Response:
 - (5)
 - 6
Passage:
|passage start|
{PASSAGE}
|passage end|
Response:
[…]
        This prompt improves the input representation by including indices as part of the input in a predefined format (parentheses).
[…]

### v2

@@
-I Wish to Have an Argument: Argumentative Reasoning in Large Language Models
+“I’d Like to Have an Argument, Please”: Argumentative Reasoning in Large Language Models
 Abstract
-We evaluate the ability of contemporary large language models (LLMs) to perform argumentative reasoning.
-We frame our experiments in terms of the argument mining (AM) and argument pair extraction (APE) tasks, and evaluate their ability to perform reasoning at increasing levels of abstraction in the input and output representations (e.g., arbitrary label sets, semantic graphs).
-We find that, although LLMs are able to match or surpass the state-of-the-art in AM and APE, their argumentative reasoning performance is very dependent on the input and output representation.
-We also find an “exemplar effect”, where too many exemplars increasingly become detrimental for task performance, and about $4-5$ being the optimal amount.
-Neither result extends to chain-of-thought (CoT) prompting: we find the exemplar effect to be nullified, and our results suggest that CoT allows for better performance under ill-conditioned problems.
-We hope that the work reported contributes to the improvement of argumentative reasoning in LLMs.
+We evaluate two large language models (LLMs) ability to perform argumentative reasoning.
+We experiment with argument mining (AM) and argument pair extraction (APE), and evaluate the LLMs’ ability to recognize arguments under progressively more abstract input and output (I/O) representations (e.g., arbitrary label sets, graphs, etc.).
+Unlike the well-known evaluation of prompt phrasings, abstraction evaluation retains the prompt’s phrasing but tests reasoning capabilities.
+We find that scoring-wise the LLMs match or surpass the SOTA in AM and APE, and under certain I/O abstractions LLMs perform well, even beating chain-of-thought–we call this symbolic prompting.
+However, statistical analysis on the LLMs outputs when subject to small, yet still human-readable, alterations in the I/O representations (e.g., asking for BIO tags as opposed to line numbers) showed that the models are not performing reasoning.
+This suggests that LLM applications to some tasks, such as data labelling and paper reviewing, must be done with care.
+keywords:
+[…]
+[…]
+[innertopmargin=-10pt, innerbottommargin=-10pt, innerleftmargin=3pt, innerrightmargin=-2pt, skipabove=0pt, skipbelow=0pt, backgroundcolor=blue!5, linecolor=black, outerlinewidth=0.01pt, font=, ]promptbox
+and
 1 Introduction
-Large language models (LLMs) such as GPT-4 (Open AI, 2023) are able to solve a significant amount of natural language processing tasks to a spectacular degree of accuracy.
-Lately, there has been a push to use these models to automate more complex tasks reliant on deeper comprehension of language, such as data labelling (Cheng, Li, and Bing, 2023; Liu et al., 2023), model evaluation (Zheng et al., 2023), and even paper reviews (Liu and Shah, 2023).
-Argumentative reasoning is central to all of these efforts.
-This type of reasoning encompasses formal and informal logic, and has as a prerequisite a deep understanding of language and the context.
-Plenty of work has been carried out to evaluate LLM performance in formal logic scenarios; but, in human-centered applications, such as chatbots, formal logic issues are overlooked in favor of defeasible reasoning (Walton, 2008).
-Hence, to accurately decide whether LLMs are reliable on more complex tasks, such as data annotation and human interaction, we must also evaluate their capabilities in the context of argumentative reasoning.
-In this paper we perform one such evaluation for GPT-3 (Brown et al., 2020) and GPT-4.
-We explore these model’s argumentative reasoning capabilities by altering the input and output representations and subsequently observing their performance in two argumentation-related tasks.
-We frame our experiments from an argument mining (AM) and argument pair extraction (APE; Cheng et al., 2020) perspective.
-AM requires the model to be able to recognize spans of a text as belonging to an argument; and APE tasks the model to match arguments from two texts as argument-rebuttal pairs.
-For the input and output representation evaluation, we test progressively more abstract reasoning settings: from the concrete setting, where the model is tasked to return matches from a passage (e.g., in AM it would return the spans that correspond to an argument); to the purely symbolic setting, where we require from the model to return the answer with arbitrary symbols (e.g., extract the spans, label them based on predefined rules, and only return the labels), or the input is in an abstract format such as abstract meaning representation (AMR; (Banarescu et al., 2013)) graphs.
-AMR graphs are known to aid in argument comparison and interpretation (Opitz et al., 2021).
-We also factor in the number of exemplars to test if the model’s reasoning capabilities depend on their context length, and chain-of-thought (CoT; (Wei et al., 2022b)) prompting, a prompting paradigm that conditions the model to work step-by-step by following a templatized process (“Let’s think about this step-by-step…”).
-It is known to provide good results in multiple reasoning tasks (Wei et al., 2022b; Kojima et al., 2022; Suzgun et al., 2023).
-The diagram below displays our evaluation.
-[…]
+Large language models (LLMs) such as GPT-4 [1] have shown to have spectacular accuracy on a variety of tasks.
+Hence, attempts have been made to automate more complex tasks reliant on argumentative reasoning, such as data labelling [2] and scientific paper reviews [3].
+Argumentative reasoning encompasses formal and informal logic, and requires a deep understanding of, and reasoning over, the pragmatic context.
+Hence, to understand the reliability of LLMs in these tasks we must also evaluate their argumentative reasoning capabilities.
+This goes beyond determining whether the model can generate relevant responses, and asks if it can robustly reason over the context and solve the task.
+We evaluate the argumentative reasoning capabilities of two LLMs, GPT-3 [4] and GPT-4 in two tasks, argument mining (AM) and argument pair extraction (APE [5]).
+[^1] Prompts, code, and outputs are in https://github.com/adewynter/argumentation-llms We do this by an abstraction evaluation: measuring progressively more abstract input and output (I/O) representations.
+Unlike prompt-phrasing evaluations, to which LLMs are known to be sensitive [6, 7], our evaluation maintains the task description untouched, and only alters the signature of the data.
+For example, adding line numbers to the input and requesting the output to be from a specific set (e.g., $\{0,1\}$) is a conceptually minor I/O representation change.
+This retains the task description, but requires some level of reasoning to return a correct and parseable response given the specified signatures.
+In this paper we then measure the LLMs’ argumentative reasoning capabilities indirectly, by testing their ability to robustly recognize arguments when altering the I/O.
 1.1 Findings
-In terms of raw scoring, we find that GPT-4 is able to reach SOTA performance in APE, and near-SOTA in AM. However, our analysis also shows the following:
+In terms of raw scoring, we find that GPT-4 is able to reach SOTA performance in APE, and near-SOTA in AM. However, our analysis shows that:
   1. 1.
-    Both LLMs fall short when the input and output representations are not carefully conditioned.
-Concrete reasoning does well, but low-abstraction (i.e., with “hints” as the input representation) symbolic approaches outperform it.
+    LLM scoring varies dramatically with the abstraction level, which suggests a lack of comprehension of the task.
+Note that across our experiments the task description remained fixed, and we only altered the I/O representation.
   2. 2.
-    The performance with respect to the number of exemplars behaves consistently regardless of setting, representation, or task.
-The LLMs have best performance when observing between 4 and 5 exemplars.
-More exemplars (up to 40 in some cases) cause a linear decrease in performance.
+    Symbolic prompting (that is, low-abstraction, hints-enabled inputs) bests other approaches, including chain-of-thought (CoT) [8].
   3. 3.
-    Performance drops noticeably, but remains above random, when the input representations are too abstract (e.g., AMR graphs).
-This does not hold for purely symbolic output representations (e.g., matrices), where the performance drops to near-zero levels.
+    CoT approaches are robust to abstraction, but its output distributions over abstractions are similar, which we attribute to its templatized nature.
   4. 4.
-    The results above do not extend to CoT.
-We find that, in line with the literature, CoT outperforms analogous non-CoT approaches.
-But we find that the distributions output by CoT are similar, which suggests that the effect of the input and output representations changes is mitigated by this technique.
-The “exemplar effect” mentioned above does not extend to CoT either.
+    LLM scores worsen with more exemplars, indicating poor inductive reasoning capabilities.
+We conclude that the LLMs are unable to reason reliably in an argumentative setting.
+2 Related Work
+We discuss evaluations of LLMs as it pertains to argumentative reasoning.
+For non-LLM-based approaches see [9]; and for a survey on reasoning in LLMs see [10].
+LLMs are relatively new, though they typically outperform non-LLM approaches, for example in AM [11] and argument evaluation [12, 13], in terms of raw scoring.
+Indeed, LLMs have been observed to generalize to (read: score well in) unseen tasks without training, thus raising the question as to whether they can reason about the prompts; or are just regurgitating their training data or returning semantically-close responses.
+To some, this generalization is an indication of emergent reasoning capabilities [14, 15]; but it has been posed that with better statistics this evidence of emergence disappears [16].
+Some tests, such as GPT-4’s own technical report [1], tout remarkable reasoning capabilities.
+There is also evidence to the contrary, e.g.
+in code generation [17], scientific questions [18], first-order logic under fictional worlds [19], and arithmetic [20]; more generally, tasks with significant reasoning depths cause LLMs to fail [21, 22].
+It has hence been suggested that LLMs do not actually reason, but rely on heuristics (e.g., semantic similarity) [23].
+Remark that these studies are limited to the prompts and versions of the models available then.
+It was suggested that LLMs are not meant for formal reasoning, and it is better to evaluate them in real-world (informal, inductive) scenarios [22]; yet GPT-3 cannot mimic human-like inductive reasoning [24], or understand the prompts [7].
+LLMs have also been found to not be competent in legal reasoning, due to their inability to make good arguments [25].
+LLMs may also retrieve dialogue acts, in line with their success as chatbots, but do not understand offers in negotiations (i.e., pragmatics) [26].
 3 Methodology
 3.1 Data
-Throughout this paper, we utilize the Review-Rebuttal Submission-v2 (RRv2) dataset from Cheng et al.
-(2020).
-It is comprised of $4,764$ pairs of review and rebuttal passages related to scientific article submissions.
-The test set size is $474$ pairs, and the data was human annotated by $5$ annotators following a specific guideline.
-Each passage is sentence-separated, and normally includes multiple arguments.
-RRv2 is a comprehensive corpus and focuses on long-distance relationships between statements; in addition to having two tasks using the same data (AM and APE), which are both actively researched in argumentative reasoning.
-For AM, each sentence is labelled with a BIO scheme, and the model must retrieve (label) each sentence from the review and rebuttal entries.
-[^1] In the RRv2 corpus, the BIO tags correspond to the Beginning, Inner, and Outer parts of an argument.
-In this task, the distinction between review and rebuttal is not relevant and each entry is treated as a separate point in the corpus.
-For APE, the arguments within each review-rebuttal pair must now be aligned: every argument made by a reviewer must be mapped, when applicable, to a refutation from the rebuttal.
-This is normally represented as a binary matrix with overlaps (Bao et al., 2022; Cheng et al., 2021).
-See Figure 1 for examples of the tasks and its labelling scheme.
+Throughout this paper, we utilize the Review-Rebuttal Submission-v2 (RRv2) dataset [5].
+It is a comprehensive corpus focused on long-distance relationships between statements, and includes both AM and APE.
+It has $4,764$ ($474$ for test) pairs of review and rebuttal passages related to scientific article submissions.
+Each passage is sentence-separated, and includes multiple arguments.
+It is human-labelled.
+For AM, each sentence is labelled with a BIO tag, [^2] In the RRv2 corpus, the BIO tags correspond to the Beginning, Inner, and Outer parts of an argument.
+and the model must retrieve (label) each sentence from the review and rebuttal entries.
+In AM the distinction between review and rebuttal is irrelevant: each entry is treated as a separate point in the corpus.
+For APE, the task is to align the arguments within each review-rebuttal pair: every argument made by a reviewer must be mapped, when applicable, to a response from the rebuttal.
+This is normally represented as a binary matrix with overlaps [27, 28].
 Prior to use we clean the text from tag and sentence delimiters.
-2.2 Models Evaluated
-We evaluate GPT-4 and the text-davinci-003 variant of GPT-3.
-The latter is similar to the original GPT-3 model, but further pretrained with reinforcement learning with human feedback (Ouyang et al., 2022; Christiano et al., 2017) and instruction pretraining (Wei et al., 2022a; Sanh et al., 2022; Ouyang et al., 2022).
-It is sometimes referred to as GPT-3.51, and has a context length of $4,097$ tokens.
-[^2] https://platform.openai.com/docs/models
-Figure 1: Examples of the (raw) data entries in RRv2.
-The BIO tag is in parentheses, and highlighted in orange and blue the “B” and “I” labels corresponding to an argument.
-In AM, review and rebuttal (response) passages are considered independent, and the task is to label all the arguments from each passage.
-In APE, the task is to match arguments from the review to their corresponding refutations.
+See Figure 1 for a sample of the corpus.
+Figure 1: Examples of the entries in RRv2.
+The BIO tag is in parenthesis.
+Highlighted in orange and blue are the “B” and “I” labels corresponding to an argument.
+In AM, review and rebuttal (response) passages are independent, and the task is to assign BIO tags to every line.
+In APE, the task is to match arguments from the review to their corresponding responses.
 In the above the first argument is unmatched, and the second (“BERT produces…”) pairs to the first argument from the response.
-Few details have been released around the architecture, model size, or pretraining procedure for GPT-4.
-Its authors consider it marginally better than GPT-3 at some tasks, and better at more complex tasks.
-They evaluate it in benchmarks specific to GPT-4 that do not explicitly include AM or APE.
-However, some of these benchmarks have an argumentation component, such as the essay portion of the SAT.
-[^3] A test for entrance to universities in the United States.
-The variant of GPT-4 we utilized has a context length of $32,768$ tokens.
-Throughout our experiments, for both models we set the temperature to $0.8$, set the maximum return tokens based on the task, and left all other parameters as their default.
-To account for randomness, we report the average of five calls.
+3.2 LLMs Evaluated
+We evaluate GPT-4 and the text-davinci-003 variant of GPT-3 (“GPT-3.51”).
+Both models are autoregressive language models, instruction-pretrained [29, 30] and tuned with reinforcement learning with human feedback [30, 31].
+For GPT-4, there are no details released around the architecture, model size, or training data.
+It is considered better than GPT-3 at more complex tasks [1].
+The variant of GPT-4 we used (“GPT-4-0613”) has a context length of $32,768$ tokens; GPT-3 has $4,097$ tokens.
+[^3] https://platform.openai.com/docs/models.
+Throughout our experiments, for both LLMs we set the temperature to $0.8$, the maximum return tokens based on the task, and left everything else as default.
+To account for randomness, we report the average of five calls per point to the Azure OpenAI API.
 3.3 Prompting
-The models are well-known to be sensitive to the prompt’s wording (Lu et al., 2022), so prior to starting the work carried out some small experimentation to find out which phrasing would be best.
-We always include exemplars of the desired output format, and, when the prompts used $n$ exemplars, we used the first $n$ points from the development set.
-In the case of AM, we also prompted GPT-4 with and without CoT, as it has been shown to improve model’s reasoning capabilities in various tasks.
-We wrote all our prompts following the structure from Open AI (2023), as we observed it produced more reliable outputs.
-For CoT we followed the standard phrasing (“Let’s read line-by-line and solve this step by step”) and indicated which sentence was being read, as well as the rationale.
+Prior to starting the work we tuned the prompt phrasing for best performance.
+When using $n$ exemplars, we used the first $n$ points from the development set.
+For AM we prompted GPT-4 with and without CoT.
+CoT conditions the LLM to work step-by-step by following a templatized process (e.g., “Let’s think about this step-by-step…”).
+It is known to provide good results in multiple reasoning tasks [8, 32, 15].
+All our prompts followed the structure from GPT-4’s technical report [1], as we observed it produced more reliable outputs.
+For CoT we followed a tuned template (“Let’s read line-by-line and solve this step by step”) indicating which sentence was being read, as well as the rationale.
 For example, “We now read {SENTENCE}.
-It follows the previous argument, and hence it is labelled with an ‘I’.” Sample prompts can be found in Figures LABEL:fig:sampleamconcrete and 2.444The full set of the prompts, scoring code, and model responses will be available at https://github.com/adewynter/argumentation-llms We used the Open AI API for all our experiments, and a consumer-grade laptop for the data processing.
-2.4 Scoring
+It follows the previous argument, and hence it is labelled with an ‘I’.” Sample prompts can be found in Figures LABEL:fig:sampleamconcrete and 2(a).
+RRv2 is measured with binary F1 (F01) for APE, and micro-F1 (Fμ) for AM [5]. Our prompts have specified a return format to signal the beginning of parsing.
+3.4 Baselines
+Our baselines are the MLMC [28] and MRC-APE [27] models.
+MLMC approaches APE as a table-filling problem: passages are related by their pairing on a table and it relies on an especially designed encoding scheme and loss.
+MRC-APE phrases it as a reading comprehension task: first, the model does AM, and then pairs the detected arguments.
+This approach is effective when using longer-context layers, of up to $4,096$ tokens.
+We additionally consider random guessers for AM (around 33% Fμ) and APE (14% F01).
 […]
-2.5 Baselines
-[…]
-[…]
-Extract from the passage all the arguments. Give
-every argument in its own paragraph.
+Extract from the passage all the arguments.
 The output should be in the form:
@@
 For example,
-{EXEMPLARS}
+{EXEMPLARS GO HERE}
 Passage:
-{PASSAGE}
+{PASSAGE GOES HERE}
 Response:
@@
 […]
-    Exemplars, if any, are included in the prompt and follow the output representation.
-    Otherwise (in the case of zero-shot), we only leave the specified output representation.
-More complex prompts, such as the one for APE, simply ask for the pairing (e.g., return all arguments from the response that match those of the review”).
-[…]
+In APE we ask for the pairing (e.g., return all arguments from the response that match those of the review”).
 […]
@@
 Extract from the passage all the arguments.
-Label the beginning of every argument with a B”.
-Label the rest of the argument with an I”.
-Label every line that is not part of an argument
-with an O”.
+Label the beginning of every argument with a
+"B". Label the rest of the argument with an
+"I". Label every line that is not part of an
+argument with an "O".
 The output should be in the form:
@@
 I
-O
 etc
@@
 For example,
-{EXEMPLARS}
+{EXEMPLARS GO HERE}
 Passage:
-{PASSAGE}
+{PASSAGE GOES HERE}
 Response:
 |begin response|
-Figure 2: Sample symbolic AM prompt.
-The model is requested to parse each line and return BIO tags.
-Exemplars, if any, are included in the prompt, and follow the specified output representation.
-For zero-shot, we only specify the output representation.
-In other output representations the model is requested to mark the B-label (e.g, “(15) 16 17”).
-In APE we request the index-based response with matching arguments (e.g., “15, 16, 17: 9, 10, 18, 19”) and convert it into the binary matrix representation for scoring.
-We also experimented with requesting the binary matrix outright.
+(a) Sample symbolic AM prompt.
+The model must return BIO tags.
+In other representations (e.g., indices) the model must to mark the B-label in parenthesis (e.g, “(15) 16 17”).
+In APE this output is of the form “argument lines: response lines”, and we convert into a binary matrix for scoring.
+Figure 2: Sample prompts for our concrete (left) and symbolic (right) settings.
+Exemplars, if any, are included in the prompt.
+For zero-shot we only specify the output representation.
+Note how the actual task definition and prompt structure remains unchanged, and we only alter the I/O representations.
 3.5 Settings
-We have named our settings (i.e., representations) as concrete and symbolic, to broadly distinguish the approach taken towards representing the task.
-The former returns full sentences, while the later encompasses a variety of input and output symbols.
-This is solely for practical purposes: in reality, symbolic approaches cover a variety of abstractions which may be considered “easier” than the concrete approach, and, strictly speaking, the concrete approach is a symbolic approach (Brachman and Levesque, 2004).
-Below we describe these settings, and we discuss our distinctions further in 3.3.
-2.6.1 Concrete
-We instruct the model to return a portion of the text based on the instructions given: in AM, the lines corresponding to an argument, and in APE, the argument pairs themselves.
+We have named our settings (i.e., representations) as concrete and symbolic, to distinguish the approach taken towards representing the task.
+Concrete returns full sentences, while symbolic encompasses a variety of I/O symbols.
+This is only for practical purposes: symbolic approaches cover multiple I/O representations, some of which may be easier than concrete; and, strictly speaking, the concrete setting is a type of symbolic representation [33].
+See Table 1 for a full description of the settings tested.
+The concrete setting we instruct the LLM to return lines in text based on the prompt: in AM, it must be part of an argument, in APE, an argument pair.
 To distinguish the “B” and “I” labels, we enforce a specific return format to work with our parsing code via a special token (|START|).
-For scoring concrete answers, we always expect an exact match.
-See Figure LABEL:fig:sampleamconcrete for a basic AM concrete prompt.
-2.6.2 Symbolic
-The model is instructed to return symbols (labels) based on its judgment of the text.
-This requires more reasoning steps than the concrete setting, given that the model is solving AM and labelling the span with a pre-specified, arbitrary label set from the instructions.
-See Figure 2 for an example.
-In the AM symbolic setting we evaluated two types of labels: the BIO tags and the line indices.
-For APE we evaluated the line indices and the full matrix representation.
-For AMR, we used amrlib [^5] https://github.com/bjascob/amrlib and a pre-trained BART-based model (Lewis et al., 2020) included in the library.
-The AMR graphs were created at the sentence-level.
+For scoring concrete settings we expect an exact text match.
+In symbolic settings the LLM must return symbols (labels) based on the prompt.
+This requires more reasoning steps than in concrete settings: the LLM is solving AM and labelling the span with an arbitrary label set defined in the prompt.
+In AM symbolic we evaluated two types of labels: BIO tags and line indices.
+For APE we evaluated line indices and the full binary matrix representation.
+We also used abstract meaning representation (AMR [34]) graphs, which are used in argument interpretation [35].
+Task Input Output (label set) Abstraction
+AM* Text Text and |START| Lowest
+AM* Text with indices Indices Low
+AM* Text Indices Medium
+AM* Text BIO tags Medium-high
+AM Text with AMR graph BIO tags High
+AM AMR graph BIO tags Highest
+Table 1: Input representations tested, in roughly increasing order of abstraction.
+Tasks marked with an asterisk (*) were tested with and without CoT.
+The first row is our concrete reasoning setting.
+Our ranking of abstraction is arbitrary: we consider the text with indices marked inline less abstract than a text without them, since the former provides a “hint” of what the label set is supposed to be like.
+Output representations with BIO tags as the output are more abstract, since it requires rule matching to determine the labeling.
 4 Experiments and Results
+We report our results comparing raw scores with respect to our settings (Section 4.1); number of exemplars (Section 4.2); and I/O representations (Section 4.3).
 4.1 AM/APE: Symbolic and Concrete Reasoning
-In our first experiment we evaluated the performance of the models in AM and APE in both the concrete setting and various iterations of symbolic settings.
-The results for the best-performing prompts and settings are in Table 1, and a description of every symbolic setting in Table 2.
-[…]
-In AM, the best-performing symbolic setting had line indices included in the input representation, and requested the indices of each argument as the output representation.
-To handle the conversion to a BIO label set, we instructed the model to return the “B” labels as indices enclosed in parentheses.
+Results for the best-performing prompts and settings are in Table 2, and a description of every setting in Table 1.
+In AM the LLMs did well but did not beat the SOTA.
+The best-performing symbolic setting had line indices included in the input representation, and requested the indices of each argument as the output representation.
+To convert to BIO tags, we instructed the model to return the “B” labels as indices enclosed in parentheses.
 Not including the indices in the input did not lead to an equivalently good performance.
-In APE, requesting a full binary matrix output led to extremely poor performance ($9.88\%F_{01}$, well below the random baseline, and closer to guessing all zeros every time).
-However, both symbolic and concrete approaches performed well with respect to the existing non-LLM-based baselines.
+For APE, GPT-4 consistently bested the best-performing models ($+14\%$ F01).
+Both symbolic and concrete approaches did well with respect to the existing non-LLM-based baselines.
 We tested other approaches, such as first extracting the arguments and then matching them, but it did not yield sufficiently good results.
+Requesting a binary matrix output led to extremely poor performance ($9.88\%$ F01, below random).
 Due to token-length and budget limitations, we were unable to test CoT and AMR in APE.
-CoT approaches had generally better performance than their non-CoT counterparts, even though they use fewer exemplars due to their higher token usage.
-The only exception to this observation was the best-performing approach (AM: indices inline and indices as output) where the difference was of $3\%$ points–still above the concrete setting.
-[…]
-GPT-3 (c) $39.86\pm 0.51$ $18.58\pm 0.70$
-GPT-4 (c) $64.51\pm 0.53$ $53.84\pm 0.73$
-GPT-3 (s) $62.00\pm 0.32$ $20.15\pm 0.91$
-GPT-4 (s) $70.63\pm 0.21$ $49.85\pm 0.96$
+CoT approaches had generally better performance than their non-CoT counterparts, even though they use fewer exemplars.
+The only exception to this was symbolic prompting (indices inline and indices as output) where the difference was $3\%$ points.
+Model AM Fμ APE F01
+GPT-3 (concrete) $39.86\pm 0.51$ $18.58\pm 0.70$
+GPT-4 (concrete) $64.51\pm 0.53$ $53.84\pm 0.73$
+GPT-3 (symbolic) $62.00\pm 0.32$ $20.15\pm 0.91$
+GPT-4 (symbolic) $70.63\pm 0.21$ $49.85\pm 0.96$
 MRC-APE $72.43$ $39.92$
 MLMC $71.35$ $32.81$
-Table 1: Results for the AM and APE tasks in the symbolic (s) and concrete (c) settings.
-We report the best-performing prompt for an average of five calls: in the case of symbolic reasoning, the best performing approach in both tasks had indices inline and indices as the output.
+Table 2: Results for the AM and APE tasks in our settings.
+The best-performing symbolic prompt had indices inline and indices as the output.
 We also report MLMC and MRC-APE, the two best-performing, non-LLM-based approaches for RRv2.
-Overall, GPT-4 had good performance, almost matching the existing baselines in AM and outperforming them in APE.
-GPT-3 had much lower performance.
+GPT-4 almost matched the existing baselines in AM and bested them in APE.
 4.2 Performance and Number of Exemplars
-On our second experiment we evaluated the performance of the LLMs in AM and APE, but compared with the number of exemplars seen by the LLMs.
-The number of exemplars was $\{0,4,8,16,\tau\}$, where $\tau$ is the maximum number allowed by the model’s token limit.
-In the case of GPT-3, $\tau$ tended to be around $4$; for GPT-4 it was a variable number depending on the task and setting, with an average of $44$ for symbolic, non-CoT approaches in AM ($15$ CoT, $9$ AMR) and $29$ concrete; and $22$ for both symbolic and concrete in APE.
-The total size of our collected data, including all exemplars, tasks, settings and five calls, is about 3.4GB.
-The results are in Figures 3 and 4.
-The models showed on average a peak in performance when the number of exemplars was $4$, and their performance steadily decreased as this number increased.
-This peak was independent of the task and the setting, since all curves were similar to one another.
-In the case of CoT, we did not observe such a trend.
-Figure 3: Effect of the number of exemplars in AM.
-Thick blue line is the concrete approach, thickest red line is the average symbolic performance.
+We compared the number of exemplars ($\{0,4,8,16,\tau\}$, where $\tau$ is the maximum number) with the LLM performance.
+For GPT-3, $\tau$ tended to be around $4$; for GPT-4 it varied, with an average of $44$ for symbolic, non-CoT approaches in AM ($15$ CoT, $9$ AMR) and $29$ concrete; and $22$ for both approaches in APE.
+Results are in Figures 3(a) and 3(b).
+The LLMs peaked in performance at $4$ exemplars, and their scores decreased from there.
+This was independent of the task and setting.
+We did not observe this trend in CoT.
+(a) Number of exemplars versus Fμ in AM. CoT (thin solid red lines) outperformed concrete and all but one of the symbolic approaches.
+(b) Number of exemplars versus F01 in APE. We could not test CoT and AMR due to limitations on token length and budget.
+Figure 3: Exemplar number versus score for AM and APE.
+In blue is the concrete approach; thick red line is the averaged performance (symbolic; dashed red lines).
 Last entry in all plots is the average maximum number of exemplars supported.
-On average, the LLMs performance peaked at $4$ exemplars, and steadily decreased as the exemplars became more numerous.
-The exception to this observation was CoT prompting (solid red lines), where they consistently outperformed concrete prompting approaches and all but one of the symbolic approaches (dashed red lines), and did not show this decreasing trend.
-GPT-3’s performance was quickly outclassed by GPT-4, even under the same number of exemplars.
-Figure 4: Effect of the number of exemplars in APE.
-Thick blue line is the concrete approach, thickest red line is the average symbolic performance.
-Last entry on each plot is the average maximum number of exemplars supported.
-Similar to AM, on average model performance peaked at $4$ exemplars, and steadily decreased as they became more numerous.
-We were unable to test CoT and AMR under this approach due to limitations on token length and budget.
-Dotted lines for symbolic are the various symbolic representations.
+LLM scoring in non-CoT peaked at around $4$ exemplars, and decreased afterwards.
 4.3 Performance and Input Representation
-On our last experiment we explored whether the representation of the input text into progressively more abstract ways led to a change of performance in the symbolic performance of the models.
-In this section we focused solely on GPT-4 and AM, since the AMR graphs tended to be too big for the models to handle (for GPT-3 in both AM and APE, for GPT-4 only APE).
-We tested three representations (text with indices inline, plain text, and with and without an AMR graph).
-There is no rigorous way to quantify the level of abstraction for these approaches.
-However, we consider the concrete approach to be least abstract, “hints” (indices inline, indices in output) to be slightly more abstract, and purely symbolic input representations (AMR graphs) as the most abstract.
-In between these categories, we ranked them based on the complexity of the output representation: BIO tags are more abstract than indices, since they have rules for matching, and both are more complex than concrete approaches since outputting the matching sentence is simpler than mapping to an arbitrary symbol and returning it.
-The full list of experiments is in Table 2, and our results in Figure 5.
-In non-CoT approaches, we observed noticeable improvements when the abstraction level was sufficiently low (indices inline and indices in output; concrete), and as the abstraction increased, the model started to perform poorly, but retained above-random performance.
-These results are significant under a Welch’s $t$-test carried out on the predictions arrays: when the input text is unaltered, requesting line numbers or BIO tags as the output provides no difference in predictions with CoT ($p\approx 0.77$; large $p$-values imply the distributions have identical expected values), but does provide a noticeable difference without ($p<0.05$).
-[…]
-Task Input Output (label set)
-AM* Text Text and |START|
-AM* Text with indices Indices
-AM* Text Indices
-AM* Text BIO tags
-AM Text with AMR graph BIO tags
-AM AMR graph BIO tags
-Table 2: Input representations tested, in roughly increasing order of abstraction.
-Tasks marked with an asterisk (*) were tested with and without CoT.
-The first row (input text, output text requested) corresponds to our concrete reasoning setting.
-Our ranking of abstraction is arbitrary: we consider the text with indices marked inline less abstract than a text without them, since the former provides a “hint” of what the label set is supposed to be like.
-Output representations with BIO tags as the output are more abstract, since it requires some rule matching in order to determine which symbol corresponds to which part of the problem.
-AMR graphs are considered the most abstract.
-Figure 5: Effect of abstraction level in the input/output representation with respect to output performance in AM.
-For non-CoT approaches, the more abstract (under our criterion) the input representation, the more difficult it is for the model to solve the task.
-For example, if the input is conditioned to have “hints” about the desired output (for example, requesting indices and adding line numbers to the input text) the performance increases significantly.
-That said, even the most abstract representations (AMR graphs) outperform noticeably the random baseline.
-CoT approaches (at the top, marked with an asterisk) consistently outperformed their non-CoT counterparts, even when the maximum number of exemplars supported is much lower, and regardless of abstraction level.
+In this section we focused on GPT-4 and AM and the following representations: text with indices inline, plain text, and with and without an AMR graph.
+There is no rigorous way to quantify the level of abstraction for these.
+However, we consider the concrete approach to be least abstract; “hints” (indices inline, indices in output) to be slightly more abstract; and purely symbolic input representations (AMR graphs) as most abstract.
+Other I/O representations are ranked based on the output representation: BIO tags are more abstract than indices (they require rules for matching); and both are more complex than concrete settings, since outputting a matching string is easier than mapping to an arbitrary symbol.
+The list of experiments is in Table 1, and our results in Figure 4.
+For non-CoT approaches, we observed noticeable improvements in low-abstraction scenarios (indices inline and indices in output; concrete).
+As it rose, LLMs scored worse, though remaining above random.
+The results are significant under a Welch’s $t$-test on the prediction arrays.
+When the input text is unaltered, line numbers or BIO tags make no difference in predictions with CoT ($p\approx 0.77$; large $p$-values imply the distributions have identical expected values), but are noticeably worse without ($p<0.05$).
+Another $t$-test shows that $t<-0.86$ and $p\approx 0.39$ for inline indices when compared with its non-CoT version.
+Hence we reject the null hypothesis that the distributions are distinct.
+Since the performance in non-CoT was better on average (67% vs 70% Fμ), it is possible that CoT harms performance on outlier points.
+We compared CoT in other scenarios, and also observed large $p$ values when comparing with only indices in the output ($<0.98$); and when comparing the latter with BIO tags ($p<0.77$).
+This suggests that CoT is effective in highly abstract scenarios, such as requesting BIO tags, but detrimental otherwise.
+Figure 4: Effect of abstraction in the I/O representation with respect to Fμ in AM.
+For non-CoT, the more abstract the input, the more difficult it is for the model to solve the task.
+Even the most abstract representations (AMR graphs) are noticeably better than a random.
+In CoT (top, marked with an asterisk) consistently outperformed non-CoT, even when the maximum number of exemplars supported is much lower, and regardless of abstraction level.
 We did not test CoT with AMR in AM due to token limitations.
 5 Discussion
-4.1 Relationship of F-scores and Reasoning
-The models, and GPT-4 in particular, performed well in both AM and APE, to the point of beating, or almost matching, the existing SOTA models.
-This alone is not sufficient to claim that the models are able to perform argumentative reasoning, let alone understand it.
-When altering the input and output representation with conceptually minor changes–such as adding line numbers or requesting a BIO set as opposed to integers–the models had noticeably different performances.
-We argue that this is indicative of a potential false positive: the models are unable to reason in an argumentative setting, but give excellent appearance of being able to do so.
-It is worth noting that, although GPT-4’s performance in APE is SOTA, it is still well-below the equivalent performance in AM.
-Although the tasks are slightly different, another plausible explanation is that the models failed to generalize due to the length of the task, which it is known to cause problems in transformer-based models (Anil et al., 2022).
-4.2 Impact of CoT in Argumentative Reasoning
-That said, we noticed that this performance difference did not extend to the inclusion of CoT.
-These prompts had on average better performance than their non-CoT counterparts, and we observed that it yielded better results in ill-posted (here, overly abstract) problems.
+The LLMs scored highly in AM and APE, to the point of beating or almost matching the existing SOTA models.
+This is not sufficient to claim that the models are able to perform argumentative reasoning.
+When we altered the I/O representation with conceptually minor changes–such as adding line numbers or requesting a BIO set as opposed to integers–the LLMs had noticeably different performances.
+An explanation for the relatively low scores in APE is that the LLMs failed to generalize due to the length of the task; which causes problems in transformer-based models [21].
+CoT prompts had on average higher scores than their non-CoT counterparts.
+They also yielded better scores in ill-posed (overly abstract) scenarios.
 However, our analysis showed that the output distributions for all CoT approaches were rather similar.
-This is perhaps indicative that CoT allows the models to return the desired output regardless of representation.
-One possible explanation for the similar output distributions is due to the templatized nature of CoT: it usually starts with a common phrase (in our case, “Let’s read line-by-line and solve this step by step”) and follows very specific steps that represent (in our prompt) the reasoning required.
-While we don’t discard this possible explanation, it is worth noting that the input and output representations are completely different in all approaches.
-That said, it is clear that CoT mitigates the effect of input and output representations on the execution of the task.
-4.3 The Exemplar Effect
-Perhaps the most surprising aspect of our evaluation was the effect of the number of exemplars with respect to the model’s downstream performance.
-It was not a linear relationship, and instead there was a clear “sweet spot” at around $4$ samples.
-Originally, Brown et al.
-(2020) had indicated that more exemplars in GPT-3 would yield better performance, but that this number must be finetuned ahead of time.
-We observed the same thing, but these results did not hold for GPT-4.
-What this suggests, from our experiment results, is that (assuming that more exemplars imply better understanding of the task) the models are not accurately performing inference from the data provided.
-That said, this “exemplar effect” did not extend to CoT settings, since the maximum number of tokens used had slightly higher performance than the other approaches.
-However, we do not discard the possibility that models with longer token limitations could also show a similar exemplar effect in CoT.
+This is perhaps indicative that CoT allows the models to return the same output regardless of representation.
+This appears to be due to the templatized nature of CoT, i.e., “Let’s read line-by-line and solve step-by-step”, and the specific steps needed to generate the output regardless of the I/O representation.
+This itself is not indicative of reasoning.
+Finally, there is a clear peak at four exemplars with respect to the model’s downstream performance.
+This suggests that, assuming that more exemplars imply better information about the task, the LLMs are not accurately performing inference from the data provided.
+This exemplar effect did not extend to CoT settings, though we do not discard the possibility that models with longer token limitations could also show this trend.
+Overall, we pose that the models are unable to reason in an argumentative setting, but their scores give an excellent appearance of being able to do so.
 6 Limitations
-5.1 Reasoning Evaluation
-It could be argued that our results are not complete in terms of evaluating argumentative reasoning capabilities.
-We agree: recognizing an argument is not the same as deciding whether it is sound, valid, or neither.
-However, without the ability of the model to show that it is able to recognize arguments and identify relations between them, any potentially generated argument or result evaluating the model’s performance in these tasks could be deemed as untrustworthy.
-Note that issues arising from the model not understanding the prompt (Webson and Pavlick, 2022) are not present in our evaluation, since the results reported here all perform better than the random baselines, and our analysis was comparative in nature.
-5.2 Data contamination
-We factored out, to an extent, potential data contamination, which is known to impact downstream model performance (Carlini et al., 2023; Liang et al., 2022; Lee et al., 2023; De Wynter et al., 2023), by tasking the model to recognize arguments from the passage.
+Our analysis has three main limitations.
+In terms of reasoning evaluation, it could be argued that our results are not complete in terms of evaluating argumentative reasoning capabilities.
+We agree: recognizing an argument is not the same as deciding its quality.
+However, without the ability of the model to show that it is able to recognize arguments and identify relations between them, any potentially generated argument or result evaluating the model’s performance in these tasks is untrustworthy.
+We factored out, to an extent, potential data contamination, which is known to impact downstream model performance [36, 37], by tasking the model to recognize arguments from the passage.
 However, we are unable to guarantee that the models have not been trained with this data, and therefore have at least some bias towards these results.
-5.3 Model Setup
-We only evaluated two models, so our results may not extend to other LLMs.
-We did not fine-tune the models, and opted instead to treating them as generalists performing in-context learning, in line with their standard, contemporary usage.
-However, it is likely that fine-tuning or specialized pre-training could show more insights into our results, and probably a more uniform performance distribution across representations: after all, MRC-APE and MLMC are specialist models and remain the SOTA in AM for RRv2.
-6 Related Work
-6.1 Informal Reasoning and Argumentation
-In this section we limit ourselves to discuss informal reasoning and argumentation evaluation of LLMs.
-For an introduction to AM covering other modelling approaches see the survey by Lawrence and Reed (2020) and for a broader introduction to the field and open problems see Lauscher et al.
-(2022).
-There has been work on using LLMs for AM (Zhao et al., 2023), as well as determining argument quality (van der Meer et al., 2022; Holtermann, Lauscher, and Ponzetto, 2022).
-More recently, Han et al.
-(2022) showed that GPT-3 is unable to mimic human-like inductive reasoning.
-While non-LLM-based approaches to APE do exist (Cheng et al., 2020; Bao et al., 2022), our work is the first to evaluate LLMs for this particular task.
-A subset of informal reasoning, commonsense reasoning, is a very active of research.
-It is not strictly the same as argumentative reasoning, but it allows for the evaluation of some types of defeasible arguments.
-While not the direct subject of this work, there are reports that GPT-3’s performance is rather low in commonsense reasoning (Cong, 2022) and figurative language (Chakrabarty, Choi, and Shwartz, 2022).
-In addition, there are techniques to improve the model’s performance in this area, such as generating explanations for out-of-distribution observations (Brahman et al., 2021), corpora and benchmarks (Forbes et al., 2020; Rudinger et al., 2020; Pyatkin et al., 2023), and decoding (Cohen and Mooney, 2023) and prompting (Liu et al., 2022; Jin et al., 2022; Štefánik and Kadlčík, 2023; Lu et al., 2022) strategies.
-Of note, Webson and Pavlick (2022) explored whether GPT-3 and other LLMs were able to understand the prompts.
-They found evidence to the contrary, and although they provided two alternate interpretations for their findings (the task was too hard or the model refused to answer), their main conclusion is that the models do not always follow, or understand, the instructions the way a human being would.
-6.2 Formal Reasoning
-The scientific community has long been interested in the evaluation of reasoning capabilities for statistical models, and LLMs are no exception ((Zhang et al., 2023); (Dziri et al., 2023); (Dan, Bastani, and Roth, 2022); (Liu et al., 2022); to name a few).
-Benchmarks like BIG-Bench Hard (Suzgun et al., 2023) have subsets of data that are meant to evaluate formal reasoning.
-Saparov and He (2023) performed a formal analysis of the reasoning capabilities of LLMs.
-They built an ontology and tested whether GPT-3 was able to perform modus ponens given a specific premise.
-They found that the models are unable to do so consistently, with a much lower performance under fictional worlds.
-This has been taken beyond formal logic and into real-world reasoning problems (e.g., maths problems), applying GPT-3, and other LLMs with much success (Open AI, 2023; Lewkowycz et al., 2022; Zong and Krishnamachari, 2023).
-That said, there exists evidence to the contrary on some tasks, such as arithmetic (Dziri et al., 2023).
-The latter reference also indicates that there is strong evidence that tasks with significant reasoning depths will cause these models to fail, a result also shared by Anil et al.
-(2022).
+Finally, we only evaluated two models, so our results may not extend to other LLMs.
+Likewise, we did not fine-tune the models, and opted instead to treating them as generalists performing in-context learning, in line with their contemporary usage.
 7 Conclusion
-In this paper we evaluated the argumentative reasoning capabilities of GPT-3 and GPT-4.
-We altered the input and output representations of the data to try to understand their behavior via their performances in AM and APE.
-We found that the models present good performance in these tasks, but this was not sufficient to conclude that they were indeed performing any type of reasoning.
-On the contrary, our experiments showed that the input and output representations had significant impact on their downstream performance.
-This sensitivity suggested that the model’s applicability to critical uses such as data labelling and paper reviewing needs to be exercised with care.
-However, when the input and output representations included CoT, our results became more surprising: the “exemplar effect” (too many exemplars being detrimental for performance) was no longer present, and the output distributions were very similar regardless of input and output representations.
-We were unable to conclude that CoT helped argumentative reasoning in LLMs, but we did note that it helped mitigate potential issues associated with ill-conditioned problems.
-As mentioned in the Limitations section, our work may be expanded by exploring the ability of the models to working in-depth with argumentation.
-Our work did not evaluate their ability to judge an argument’s strength, or for the model to provide reasonable rebuttals.
+We evaluated the argumentative reasoning capabilities of GPT-3 and GPT-4, by measuring whether they could recognize arguments from a passage–the first step on performing such reasoning.
+The LLMs score well in AM and APE, beating or nearly-matching the SOTA.
+However, statistical analysis on the LLMs’ predictions when subject to small, yet still human-readable, alterations in the I/O representations showed that the LLMs were extremely sensitive to the abstraction level and the number of exemplars.
+Hence, we concluded that they were not reasoning over the arguments seen.
+However, symbolic prompting strategies (e.g.
+reducing the abstraction level of the prompt by adding line numbers) allowed the LLMs to score well and even beat CoT.
+We were also unable to conclude that CoT helped argumentative reasoning in LLMs, but did observe more robust results due to its templatized nature.
+Our analysis implies that it helps mitigate issues stemming from overly abstract or ill-conditioned problems.
+As mentioned in Section 6, we did not evaluate the LLMs’ ability to judge an argument’s strength, or to provide reasonable rebuttals.
 Moreover, due to token length limitations, we were unable to evaluate AMR, the most abstract setting we tested, with CoT.
-We believe that the performance difference between non-CoT AMR and CoT AMR would provide valuable insights on to what extent these models are able to discern abstract input representations successfully.
-All of these are well-known problems to the computational argumentation field and the scientific community at large.
-We plan to explore in further work these as they relate to LLMs (GPT-4 specifically, given its performance compared to GPT-3) as in-depth evaluation of these is important for downstream applications.
-Finally, although the goal of our work is to study LLM argumentative reasoning capabilities, we believe that this type of representation-based analysis could be extended to other models and non-argument-related tasks, such as natural language inference.
+We believe that this evaluation could provide valuable insights on to what extent these models are able to discern abstract input representations.
+Further work could explore these issues.
+Overall, our work suggests that LLM usage in areas like data labelling and paper reviewing must be exercised with care and good judgement.
 Acknowledgements
-The authors wish to thank Liying Cheng for answering questions about the RRv2 dataset and its methodologies.
+The authors wish to thank Liying Cheng for answering questions about the RRv2 dataset.
 Appendix A Full Prompts
@@
 {EXEMPLAR REVIEW 1 GOES HERE}
- |end of review|
+|end of review|
 |start of response|
@@
 {REVIEW}
- |end of review|
+|end of review|
 |start of response|
@@
 So the answer should be:
- - O
- - O
- - O
- - O
- - O
- - B
- - I
+- O
+- O
+- O
+- O
+- O
+- B
+- I
 Passage:
@@
 Response:
- - O
- - O
- - O
- - O
- - O
- - B
- - I
+- O
+- O
+- O
+- O
+- O
+- B
+- I
 Passage:
@@
 {EXEMPLAR REVIEW 1 GOES HERE}
- |end of review|
+|end of review|
 |start of response|
@@
 - argument 0
- - I am not at all familiar with quantization methods
+- I am not at all familiar with quantization methods
 therefore no knowledge of relevant related works.
-  - If, however, the authors did a thorough job of
+- If, however, the authors did a thorough job of
 surveying related works and chose sensible baselines
@@
 - rebuttal 0
- - We have updated our manuscript so that it states
+- We have updated our manuscript so that it states
 our contributions explicitly.
-  - While our proposed DualPrecision is simple yet
+- While our proposed DualPrecision is simple yet
 effective, we would like to emphasize that it is
 only one of our contributions.
-  - For the reviewers convenience, we re-iterate our
+- For the reviewers convenience, we re-iterate our
 main contributions below:
-  - Overall, we systematically analyze the model size
+- Overall, we systematically analyze the model size
 and accuracy trade-offs considering both weight
@@
 Response:
- - (5)
- - 6
+- (5)
+- 6
 Passage: